Find an equation of the line that satisfies the given conditions.
Through (1/2, - 2/3); perpendicular to the line 4x − 8y = 1

Answer by Cromlix(4381)

hi there,
Sort line equation 4x - 8y = 1
into y = mx + c form
-8y = - 4x + 1
Divide throughout by -8
y = 1/2x - 1/8
Lines that are perpendicular to
one another have gradients that
multiply together to give -1
m1 x m2 = -1
In this case:
1/2 x m2 = -1
m2 = -2
Setting up the line equation
y - b = m(x - a)
Where m = -2 and (a,b) -> (1/2, -2/3)
y -(-2/3) = -2(x - 1/2)
y + 2/3 = -2x + 1
y = -2x + 1 - 2/3
y = -2x + 1/3
or multiply through by 3
3y = -6x + 1
Hope this helps :-)